White light ranging from blue (400 nm) to red (700 nm) illuminates a diffraction grating with 5000lines/cm .

What are the angles of the first-order maximum for blue and red?

?R,?B = degrees
Answer #1

for this first we calculate the part (b) andfirnd the the angles of the first-order maximum for blue

   and red

   out of these two angles which angle is low thatcolour will be closer to the central maximum

   so we get

   (e + d) sin? = n ?

   n = 1

   (e + d) = (1 / 8000)

   substituting each wave length find theangles
Answer #2

White light ranging from blue (400 nm) to red (700 nm) illuminates a diffraction grating with 6000lines/cm .

What are the angles of the first-order maximum for blue and red?

\par Angle of first-order maximum is given as \par \theta=\frac{\lambda}{d} \par Here \lambda_B=400 nm =400*10^{-9} m\par \lambda_R=700 nm =700*10^{-9} m \par_, d=\frac{1}{6000}*10^{-2} m \par Therefore for blue \par \theta_B=\frac{\lambda_B}{d}=\frac{400*10^{-9}*6000}{10^{-2}} = 0.24 ^0 \par \theta_R=\frac{\lambda_R}{d}=\frac{700*10^{-9}*6000}{10^{-2}} = 0.42 ^0